OpenStudy (anonymous):

(4x-1)^(1/2) - 1/3(4x-1)^3/2

OpenStudy (anonymous):

if there's a fraction in the exponent, take to the power of the numerator and take the root of the denominator. in the first part of this problem: \[(4x-1)^{\frac{ 1 }{ 2 }}=\sqrt{(4x-1)^{1}}\]

OpenStudy (anonymous):

would i do that for both parts? or do i also have to distribute the denominator of 3 to all sides

OpenStudy (amistre64):

im not too sure what the question is ..

OpenStudy (anonymous):

I have to factor completely this equation

OpenStudy (anonymous):

sqrt(4x-1)-1/3(4x-1)(sqrt(4x-1)=sqrt(4x-1)(1-4/3x+1) =sqrt(4x-1)(-4/3x+2)

OpenStudy (amistre64):

the greatest common factor is the: ^(1/2)

OpenStudy (amistre64):

for example: a - 4a^2 b, the gcf is a a(1-4ab)